# Cyclotomic generating functions

### v2

@@ 6.2. Gale order and the associated CGF monoid
The properties in 6.2 also hold for the Gale monoid. In particular, it has a finite number of elements of each degree and a unique minimal set of generators which can be explored computationally. Data is given in Section 7 for the number of elements in $\Phi^{\operatorname{Gale}}$ of degree $n$ up to $n=18$ along with the pointer to the corresponding OEIS entry. The number of generators of each degree are also noted in Section 7.
As noted in 3.11, not all basic CGF’s are in the Gale monoid. They agree up to degree 10, there are two basic CGF’s which don’t satisfy the Gale property in degree 11, namely
@@
There are 4 non-Gale basic CGF’s in degree 12, and so on.
The Gale monoid and unimodal CGF monoids are not comparable. The Gale monoid includes $[4]_{q}/[2]_{q}=\Phi_{4}=1+q^{2}$, which is not unimodal. The smallest degree unimodal CGF’s which don’t satisfy the Gale property have degree 20,
@@
6.3. CGF monoid from regular sequences
@@
Fix a pair of multisets of positive integers of the same size, say $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{m}\}$. Let $\mathbbm{k}[x_{1},\ldots,x_{m}]$ be a (free) polynomial ring over a field $\mathbbm{k}$ with grading determined by $\deg(x_{j})=b_{j}$ for $1\leq j\leq m$. Let $\theta_{1},\ldots,\theta_{m}$ be a sequence of non-constant homogeneous polynomials in $\mathbbm{k}[x_{1},\ldots,x_{m}]$ with $\deg(\theta_{j})=a_{j}$. Then, $\theta_{1},\ldots,\theta_{m}$ is a regular sequence if $\theta_{i}$ is not a zero-divisor in $\mathbbm{k}[x_{1},\ldots,x_{m}]/(\theta_{1},\ldots,\theta_{i-1})$ for all $1\leq i\leq m$. Furthermore, $\theta_{1},\ldots,\theta_{m}\in\mathbbm{k}[x_{1},\ldots,x_{n}]$ form a homogeneous system of parameters (HSOP) if $\mathbbm{k}[x_{1},\ldots,x_{m}]$ is a finitely generated $\mathbbm{k}[\theta_{1},\ldots,\theta_{m}]$-module. Consider the corresponding quotient rings,
@@
Lemma 6.8.
Let $\mathbbm{k}[x_{1},\ldots,x_{m}]$ be a polynomial ring over a field $\mathbbm{k}$ with $\deg(x_{j})=b_{j}\in\mathbb{Z}_{\geq 1}$. Suppose $\theta_{1},\ldots,\theta_{m}$ are homogeneous elements. Then the following are equivalent:
@@
    for all $1\leq i\leq m$, there is some $N$ such that $x_{i}^{N}\in(\theta_{1},\ldots,\theta_{m})$;
@@
    the only common zero of $\theta_{1},\ldots,\theta_{m}$ over the algebraic closure $\overline{\mathbbm{k}}$ is the origin $(0^{m})$.
@@
$\operatorname{Hilb}(R;q)\coloneqq\sum_{k\geq 0}(\dim R_{k})q^{k}$
@@
Theorem 6.9 (c.f. [Sta78, Cor. 3.2-3.3]).
If $\theta_{1},\ldots,\theta_{m}$ is a regular sequence with $\deg(\theta_{j})=a_{j}$ in the polynomial ring $\mathbbm{k}[x_{1},\ldots,x_{m}]$ with $\deg(x_{j})=b_{j}$, then $R=\mathbbm{k}[x_{1},\ldots,x_{m}]/(\theta_{1},\ldots,\theta_{m})$ is a finite-dimensional $\mathbbm{k}$-vector space, and its Hilbert series is the basic cyclotomic generating function
(6.1) $\operatorname{Hilb}(R;q)\coloneqq\sum_{k\geq 0}(\dim R_{k})q^{k}=\prod_{k=1}^{m}\frac{[a_{k}]_{q}}{[b_{k}]_{q}}\in\Phi^{+}.$
@@
(6.2) $\frac{\mathbb{C}[x_{1},\ldots,x_{k}]^{S_{k}}}{\langle h_{\ell+1},\ldots,h_{\ell+k}\rangle}$
@@
CGF’s are often intimately related to the cyclic sieving phenomenon [RSW04], and Douvropolous proved a CSP in this context [Dou18].
Suppose two basic CGF’s $f,g$ arise as the Hilbert series of quotient rings $R_{1}$ and $R_{2}$ coming from homogeneous systems of parameters. Then $fg$ arises in the same way by taking the tensor product of $R_{1}$ and $R_{2}$. Hence we may consider the HSOP monoid $\Phi^{\operatorname{HSOP}}$ as a submonoid of $\Phi^{+}$ consisting of all Hilbert series of quotients $\mathbb{C}[x_{1},\ldots,x_{m}]/(\theta_{1},\ldots,\theta_{m})$ for homogeneous systems of parameters.
Lemma 6.12.
@@
Suppose $\theta_{1},\ldots,\theta_{m}$ is a regular sequence for $\mathbbm{k}[x_{1},\ldots,x_{m}]$ where $\deg(\theta_{i})=a_{i}$, $\deg(x_{i})=b_{i}$, and we have sorted the elements so that $a_{1}\leq\cdots\leq a_{m}$ and $b_{1}\leq\cdots\leq b_{m}$. If $a_{k}<b_{k}$, then $\theta_{1},\ldots,\theta_{k}$ have degree at most $a_{k}$, and $\deg(x_{k})=b_{k}>a_{k}$, so $\theta_{1},\ldots,\theta_{k}\in\mathbbm{k}[x_{1},\ldots,x_{k-1}]$. By 6.9, $\mathbbm{k}[x_{1},\ldots,x_{k-1}]/(\theta_{1},\ldots,\theta_{k-1})$ is finite-dimensional, so it contains a power of $\theta_{k}$, contradicting the zero-divisor condition. Hence, $a_{k}\geq b_{k}$ for all $k=1,\ldots,m$ and the result follows from the definition of Gale order. ∎
Which basic CGF’s can be realized as the Hilbert series of a quotient ring by a homogeneous sequence of parameters? In the special case $b_{1}=\cdots=b_{m}=1$, i.e. $\deg(x_{j})=1$ for all $j$, a homogeneous system of parameters corresponds to a sequence of generators for a classical complete intersection $X$. The corresponding cyclotomic generating function is the Hilbert series of the projective coordinate ring of $X$,
@@
For general denominator multisets $\{b_{1},\ldots,b_{m}\}$, such homogeneous systems of parameters yield complete intersections inside weighted projective space $\mathbb{P}^{\{b_{1},\ldots,b_{m}\}}$. However, it is not at all clear which multisets $\{a_{1},\ldots,a_{m}\}$ are realizable degrees for some homogeneous system of parameters. The requirement that $\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\mathbb{Z}_{\geq 0}[q]$ is a significant hurdle, as the next example shows.
Example 6.13.
@@
is not a cyclotomic generating function since it has a negative coefficient. Therefore, $\mathbbm{k}[x_{1},x_{2},x_{3}]$ with $\deg(x_{1})=2,\deg(x_{2})=3,\deg(x_{3})=7$ has no homogeneous system of parameters with degrees $3,5,14$. Indeed, in this case, the only monic homogeneous elements of degree $3$ and $5$ are $x_{2}$ and $x_{1}x_{2}$, and one may see in a variety of ways (regular sequences, dimension counting, non-trivial vanishing) that these two elements cannot belong to a homogeneous system of parameters. This example also satisfies all of the necessary conditions in 3.9.
Example 6.14.
@@
for all $c_{1},\ldots,c_{m},b_{1},\ldots,b_{m}\in\mathbb{Z}_{\geq 1}$. Therefore, every sequence of positive integers $b_{1},\ldots,b_{k}$ gives rise to a finite dimensional quotient ring with Hilbert series given by a CGF with denominator given by the product of $[b_{i}]_{q}$’s.
By 3.9, we know that for any cyclotomic generating function $f(q)=\prod_{j=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\Phi^{+}$, including $f\in\Phi^{\operatorname{HSOP}}$, we have $a_{k}\in\operatorname{Span}_{\mathbb{Z}_{\geq 0}}\{b_{1},\ldots,b_{m}\}$. It is not immediately clear how to characterize which sequences $a_{1},\ldots,a_{k}$ can actually be realized as degree sequences of HSOP’s.
Example 6.15.
@@
One way to explore the HSOP monoid with a fixed denominator multiset comes from an analogue of 3.14. In particular, the analogous HSOP subgraph of the graph described in Section 3.5 remains connected with diameter at most $2m$.
Lemma 6.19.
@@
[…]
are regular sequences for all $1\leq i\leq m$.
Remark 6.20.
The argument in 6.9 works more generally when $\mathbbm{k}[x_{1},\ldots,x_{m}]$ is a Cohen–Macaulay $\mathbb{N}$-graded $\mathbbm{k}$-algebra of Krull dimension $m$ in the sense of [Sta79, §3], except that the stated expression for $\operatorname{Hilb}(R;q)$ will be multiplied by some polynomial $P(q)\in\mathbb{Z}[q]$. Intuitively, $P(q)$ arises from syzygies of $x_{1},\ldots,x_{m}$ and may have negative coefficients.
The study of Cohen–Macaulay rings has been a rich subject. It would be interesting to consider properties of the set of all Hilbert series of Cohen–Macaulay rings as we have done here for CGF’s.
Open Problem 6.21.

### v3

@@ 6.2 Gale order and the associated CGF monoid
The properties in Lemma 46 also hold for the Gale monoid. In particular, it has a finite number of elements of each degree and a unique minimal set of generators which can be explored computationally. Data is given in Section 7 for the number of elements in $\Phi^{\operatorname{Gale}}$ of degree $n$ up to $n=18$ along with the pointer to the corresponding OEIS entry. The number of generators of each degree are also noted in Section 7.
As noted in Remark 32, not all basic CGFs are in the Gale monoid. They agree up to degree 10, there are two basic CGFs which don’t satisfy the Gale property in degree 11, namely
@@
There are 4 non-Gale basic CGFs in degree 12, and so on.
The Gale monoid and unimodal CGF monoids are not comparable. The Gale monoid includes $[4]_{q}/[2]_{q}=\Phi_{4}=1+q^{2}$, which is not unimodal. The smallest degree unimodal CGFs which don’t satisfy the Gale property have degree 20,
@@
6.3 CGF monoid from regular sequences
@@
Fix a pair of multisets of positive integers of the same size, say $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{m}\}$. Let $\mathbbm{k}[x_{1},\ldots,x_{m}]$ be a (free) polynomial ring over a field $\mathbbm{k}$ with grading determined by $\deg(x_{j})=b_{j}$ for $1\leqslant j\leqslant m$. Let $\theta_{1},\ldots,\theta_{m}$ be a sequence of non-constant homogeneous polynomials in $\mathbbm{k}[x_{1},\ldots,x_{m}]$ with $\deg(\theta_{j})=a_{j}$. Then, $\theta_{1},\ldots,\theta_{m}$ is a regular sequence if $\theta_{i}$ is not a zero-divisor in $\mathbbm{k}[x_{1},\ldots,x_{m}]/(\theta_{1},\ldots,\theta_{i-1})$ for all $1\leqslant i\leqslant m$. Furthermore, $\theta_{1},\ldots,\theta_{m}\in\mathbbm{k}[x_{1},\ldots,x_{n}]$ form a homogeneous system of parameters (HSOP) if $\mathbbm{k}[x_{1},\ldots,x_{m}]$ is a finitely generated $\mathbbm{k}[\theta_{1},\ldots,\theta_{m}]$-module. Consider the corresponding quotient rings,
@@
Lemma 52.
Let $\mathbbm{k}[x_{1},\ldots,x_{m}]$ be a polynomial ring over a field $\mathbbm{k}$ with $\deg(x_{j})=b_{j}\in\mathbb{Z}_{\geqslant 1}$. Suppose $\theta_{1},\ldots,\theta_{m}$ are homogeneous elements. Then the following are equivalent:
@@
    for all $1\leqslant i\leqslant m$, there is some $N$ such that $x_{i}^{N}\in(\theta_{1},\ldots,\theta_{m})$;
@@
    the only common zero of $\theta_{1},\ldots,\theta_{m}$ over the algebraic closure $\overline{\mathbbm{k}}$ is the origin.
@@
$\operatorname{Hilb}(R;q)\coloneqq\sum_{k\geqslant 0}(\dim R_{k})q^{k}$
@@
Theorem 53 (c.f. [Sta78, Cor. 3.2-3.3]).
If $\theta_{1},\ldots,\theta_{m}$ is a regular sequence with $\deg(\theta_{j})=a_{j}$ in the polynomial ring $\mathbbm{k}[x_{1},\ldots,x_{m}]$ with $\deg(x_{j})=b_{j}$, then
$R=\mathbbm{k}[x_{1},\ldots,x_{m}]/(\theta_{1},\ldots,\theta_{m})$
is a finite-dimensional $\mathbbm{k}$-vector space, and its Hilbert series is the basic cyclotomic generating function
$\operatorname{Hilb}(R;q)\coloneqq\sum_{k\geqslant 0}(\dim R_{k})q^{k}=\prod_{k=1}^{m}\frac{[a_{k}]_{q}}{[b_{k}]_{q}}\in\Phi^{+}.$ (6.1)
@@
$\frac{\mathbb{C}[x_{1},\ldots,x_{k}]^{S_{k}}}{\langle h_{\ell+1},\ldots,h_{\ell+k}\rangle}$ (6.2)
@@
CGFs are often intimately related to the cyclic sieving phenomenon [RSW04], and Douvropolous proved a CSP in this context [Dou18].
Suppose two basic CGFs $f,g$ arise as the Hilbert series of quotient rings $R_{1}$ and $R_{2}$ coming from homogeneous systems of parameters. Then $fg$ arises in the same way by taking the tensor product of $R_{1}$ and $R_{2}$. Hence we may consider the HSOP monoid $\Phi^{\operatorname{HSOP}}$ as a submonoid of $\Phi^{+}$ consisting of all Hilbert series of quotients $\mathbb{C}[x_{1},\ldots,x_{m}]/(\theta_{1},\ldots,\theta_{m})$ for homogeneous systems of parameters.
Lemma 56.
@@
Suppose $\theta_{1},\ldots,\theta_{m}$ is a regular sequence for $\mathbbm{k}[x_{1},\ldots,x_{m}]$ where $\deg(\theta_{i})=a_{i}$, $\deg(x_{i})=b_{i}$, and we have sorted the elements so that $a_{1}\leqslant\cdots\leqslant a_{m}$ and $b_{1}\leqslant\cdots\leqslant b_{m}$. If $a_{k}<b_{k}$, then $\theta_{1},\ldots,\theta_{k}$ have degree at most $a_{k}$, and $\deg(x_{k})=b_{k}>a_{k}$, so $\theta_{1},\ldots,\theta_{k}\in\mathbbm{k}[x_{1},\ldots,x_{k-1}]$. By Theorem 53, $\mathbbm{k}[x_{1},\ldots,x_{k-1}]/(\theta_{1},\ldots,\theta_{k-1})$ is finite-dimensional, so it contains a power of $\theta_{k}$, contradicting the zero-divisor condition. Hence, $a_{k}\geqslant b_{k}$ for all $k=1,\ldots,m$ and the result follows from the definition of Gale order. ∎
Which basic CGFs can be realized as the Hilbert series of a quotient ring by a homogeneous sequence of parameters? In the special case $b_{1}=\cdots=b_{m}=1$, i.e. $\deg(x_{j})=1$ for all $j$, a homogeneous system of parameters corresponds to a sequence of generators for a classical complete intersection $X$. The corresponding cyclotomic generating function is the Hilbert series of the projective coordinate ring of $X$,
@@
For general denominator multisets $\{b_{1},\ldots,b_{m}\}$, such homogeneous systems of parameters yield complete intersections inside weighted projective space $\mathbb{P}^{\{b_{1},\ldots,b_{m}\}}$. However, it is not at all clear which multisets $\{a_{1},\ldots,a_{m}\}$ are realizable degrees for some homogeneous system of parameters. The requirement that $\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\mathbb{Z}_{\geqslant 0}[q]$ is a significant hurdle, as the next example shows.
Example 57.
@@
is not a cyclotomic generating function since it has a negative coefficient. Therefore, $\mathbbm{k}[x_{1},x_{2},x_{3}]$ with $\deg(x_{1})=2,\deg(x_{2})=3,\deg(x_{3})=7$ has no homogeneous system of parameters with degrees $3,5,14$. Indeed, in this case, the only monic homogeneous elements of degree $3$ and $5$ are $x_{2}$ and $x_{1}x_{2}$, and one may see in a variety of ways (regular sequences, dimension counting, nontrivial vanishing) that these two elements cannot belong to a homogeneous system of parameters. This example also satisfies all of the necessary conditions in Lemma 30.
Example 58.
@@
for all $c_{1},\ldots,c_{m},b_{1},\ldots,b_{m}\in\mathbb{Z}_{\geqslant 1}$. Therefore, every sequence of positive integers $b_{1},\ldots,b_{k}$ gives rise to a finite dimensional quotient ring with Hilbert series given by a CGF with denominator given by the product of $[b_{i}]_{q}$’s.
By Lemma 30, we know that for any cyclotomic generating function
$f(q)=\prod_{j=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}\in\Phi^{+},$
including $f\in\Phi^{\operatorname{HSOP}}$, we have $a_{k}\in\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}\{b_{1},\ldots,b_{m}\}$. It is not immediately clear how to characterize which sequences $a_{1},\ldots,a_{k}$ can actually be realized as degree sequences of HSOP’s.
Example 59.
@@
One way to explore the HSOP monoid with a fixed denominator multiset comes from an analogue of Lemma 35. In particular, the analogous HSOP subgraph of the graph described in Section 3.5 remains connected with diameter at most $2m$.
Lemma 63.
@@
$\gamma_{1},\ldots,\gamma_{i},\theta_{i+1},\ldots,\theta_{m}\qquad\text{and}\qquad\gamma_{1},\ldots,\gamma_{i},\theta_{i+1}^{\prime},\ldots,\theta_{m}^{\prime}$
are regular sequences for all $1\leqslant i\leqslant m$.
Remark 64.
The argument in Theorem 53 works more generally when $\mathbbm{k}[x_{1},\ldots,x_{m}]$ is a Cohen–Macaulay $\mathbb{N}$-graded $\mathbbm{k}$-algebra of Krull dimension $m$ in the sense of [Sta79, §3], except that the stated expression for $\operatorname{Hilb}(R;q)$ will be multiplied by some polynomial $P(q)\in\mathbb{Z}[q]$. Intuitively, $P(q)$ arises from syzygies of $x_{1},\ldots,x_{m}$ and may have negative coefficients.
The study of Cohen–Macaulay rings has been a rich subject. It would be interesting to consider properties of the set of all Hilbert series of Cohen–Macaulay rings as we have done here for CGFs.
Problem 65.
@@ Acknowledgments
Billey was partially supported by the Washington Research Foundation and NSF DMS-1764012. Swanson was partially supported by NSF DMS-2348843.